# Curvature corrections to Starobinsky inflation can explain the ACT results

### v2

@@ III.1 $R^{3}$ correction
Though a series expansion in $\delta_{3}$ of the slow-roll parameters (15) is useful for understanding the qualitative behaviour of $n_{s}$ and $r$ for small $\delta_{3}$, in a more reasonable approach $\delta_{3}$ takes finite values and may not be much smaller than $y^{2}$ during inflation. This means that the approximation of the slow-roll parameters in Eq. (15) may not lead to accurate predictions for the observables, and we have to take into account all orders in $\delta_{3}$ when deriving $n_{s}$ and $r$ at the end of this Section. The same applies to the parameters at the $R^{4}$ and $R^{3/2}$ terms described below. The scalar potential $V(\phi)$ and $n_{s}-r$ plots for the three models are shown at the end of this Section, see Figs. 1 and 2.
@@ III.4 The results for $n_{s}$, $r$ and $\alpha_{s}$
These results appear to favor $R^{3},R^{4}$ models which can have $1\sigma$ compatibility with the new data, while $R^{3/2}$ can be only compatible within $2\sigma$ if inflation lasts no more than around $60$ e-folds.

### v3

@@ III.1 $R^{3}$ correction
Though a series expansion in $\delta_{3}$ of the slow-roll parameters (15) is useful for understanding the qualitative behaviour of $n_{s}$ and $r$ for small $\delta_{3}$, in a more reasonable approach $\delta_{3}$ takes finite values and may not be much smaller than $y^{2}$ during inflation. This means that the approximation of the slow-roll parameters in Eq. (15) may not lead to accurate predictions for the observables, and we have to take into account all orders in $\delta_{3}$ when deriving $n_{s}$ and $r$ at the end of this Section. The same applies to the parameters at the $R^{4}$ and $R^{3/2}$ terms described below. The scalar potential $V(\phi)$ and $n_{s}-r$ plots for the three models are shown at the end of this Section, see Figs. 1 and 2. Figure 1 shows that the potentials for the $R^{3}$-model (as well as the $R^{4}$-model discussed below) become steep close to the discontinuity value of $\phi$ (for negative $\delta_{3}$ and $\delta_{4}$). Having obtained numerical solutions to the inflaton equation of motion with initial conditions in the steep region at the beginning of inflation, we verified that there is no violation of slow-roll during the last 506050-60 e-folds of inflation because an inflaton velocity is suppressed by Hubble friction.
@@ III.4 The results for $n_{s}$, $r$ and $\alpha_{s}$
These results appear to favor $R^{3},R^{4}$ models which can have $1\sigma$ compatibility with the new data, while $R^{3/2}$ can be compatible within only $2\sigma$ if inflation lasts no more than around $60$ e-folds.